OpenStudy (anonymous):

Convert this percent to a decimal. 33 ------ 50 × 1% = _____

OpenStudy (anonymous):

(33*1)/(50*100) = (33/5)*(1/1000) = (3.3*10*2/10)*10^-3 = 6.6*10^-3 or \[\left( \frac{ 33 }{ 50 } \right)\times \left( \frac{ 1 }{ 100 } \right)=\left( \frac{ 3.3\times2 }{ 5\times2 } \right)\times10^{-2}=6.6\times 10^{-3}\]